OpenStudy (anonymous):

solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2 + 10x = 18 –11.56, 1.56 11.56, 1.56 –11.56, –1.56 11.56, –1.56

OpenStudy (dtan5457):

use (b/2)^2 to get your new value of C, and add to both sides. (25 in this case) x^2+10x+25=43 change to: (x+5)^2=43 take square roots \[x+5=\pm \sqrt{43}\] subtract the 5. \[x=-5\pm \sqrt{43}\]

OpenStudy (anonymous):

is it b?

OpenStudy (dtan5457):

nope